i) Kishan is a sales executive who has to make visits to Colombo, Kandy, Galle,


Gampaha, Panadura and Negombo. He has to visit each city once only and always


takes the unique shortest route between any two cities.


Kishan's manager calls and now tells him he has to start his journey in Galle


and must visit four cities in total. How many possible routes are there for


Kishan in this case? ( 4 marks)


f. If Kishan was to pick a route for the shortened four city journey at random,


what is the probability that the route he selected would be one starting from


Galle?


1
Expert's answer
2021-11-22T19:08:10-0500

i) The order does not matter


"C(5, 3)=\dbinom{5}{3}=\dfrac{5!}{3!(5-3)!}=\dfrac{5(4)}{1(2)}=10"

f.

"p=\dfrac{1}{6}"